# Consensus in Motion: A Case of Dynamic Rationality of Sequential Learning in Probability Aggregation

### v1

@@ 1 Introduction
In contrast, this paper advocates a shift toward a dynamic theory of collective decision-making in which both individual and collective beliefs update rationally in response to new information (see [5]). Motivated by the first impossibility result for dynamically rational judgment aggregation [10], our main contribution is a possibility result for dynamically rational collective decision-making on logically interconnected propositions. We introduce a framework of probability aggregation based on propositional probability logic that generalizes traditional approaches from both judgment and preference aggregation. Building on prior work in probability aggregation [17, 1, 2, 12, 18, 7, 9], our first theorem establishes that any consensus-compatible and independent aggregation rule on a non-nested agenda must be linear. Furthermore, we show that dynamically rational linear aggregation is attainable when the domain of the aggregation function is restricted to profiles in which all group members share the same probabilistic judgments for a specified subset of propositions—what we term the common ground (here understood in line with a broader notion of consensus) [^1] Throughout this paper, “consensus” is used in two senses: (i) as a broad notion of mutual agreement—here, synonymous with our technical term “common ground” and (ii) as “consensus compatibility,” the requirement that if all individuals assign probability 1 to a proposition, then the collective judgment must also assign 1.. This common ground represents the set of propositions on which all individuals already agree, and by ensuring that new information is limited to this shared foundation—a process we call fair learning—all members benefit equally, preventing dynamically irrational outcomes arising from unequal starting beliefs. Importantly, we demonstrate that this fair learning process not only achieves dynamic rationality but also preserves the common ground across sequential updates.
@@ 2.1 Propositional probability logic
Let $\mathcal{L}$ be the set of all logically non-equivalent [^2] Logical equivalence (i.e., $\phi\equiv\psi$ if and only if $\models\phi\leftrightarrow\psi$) is an equivalence relation. We select one finitary representative from each equivalence class so that no two distinct formulas in $\mathcal{L}$ are equivalent. Since $A$ is finite, and we consider only finitary formulas, $\mathcal{L}$ is finite. formulas based on a propositional logic with a finite set $A$ of atomic formulas. The variable $\phi$ is then a generic variable ranging over all well-formed formulas in $\mathcal{L}$, i.e., any expression that can be constructed from the atomic elements in $A$ using the logical connectives ¬,,,$\neg$,$\land$,$\lor$,$\rightarrow$,$\leftrightarrow$ , ∧ , ∨ , → , ↔.
@@
  • If $\phi_{1},\phi_{2}...\in\mathcal{L}$ such that $\models\neg(\phi_{i}\land\phi_{j})$ for all $i\neq j\in\mathbb{N}$, then
    $P(\bigvee_{i=1}^{\infty}\phi_{i})=\sum_{i=1}^{\infty}P(\phi_{i})$
@@ 2.2 The framework of probability aggregation
The probabilistic judgement $J$ is a function $J\colon X\rightarrow[0,1]$. The judgement $J$ is called probabilistically rational if there exists a propositional probability $P^{*}\colon\mathcal{L}\rightarrow[0,1]$ on the whole set $\mathcal{L}$ such that $\forall\ \phi\in X\colon J(\phi)=P^{*}(\phi)$. A probabilistic profile on an agenda $X$ is a tuple $(J_{1},...,J_{n})$, where each $J_{1},...,J_{n}$ is a probabilistically rational probabilistic judgement. The probabilistic aggregation rule is a function $F\colon J(X)^{n}\rightarrow J(X),$ where $J(X)$ is the collective set of all probabilistically rational probabilistic judgements on the agenda $X$.
@@ 2.3 A Running Example: A Political Scenario
Population aging is straining the healthcare and pension systems.
@@
The current tax revenue is sufficient to meet the fiscal demands imposed by population aging.
@@
Table 1: A sample probabilistic profile on the agenda $X=\pm\{a,b,c,a\rightarrow b,a\land b,a\land c,b\land c,a\land b\land c\}$, where each individual assigns a degree of confidence (ranging from 0 to 1) to the propositions. Each entry is chosen so as to align with the relationships dictated by logical connectives, ensuring that bounds for probability assignments are satisfied.
Individual $a$ $b$ $c$ $a\rightarrow b$ $a\land b$ $a\land c$ $b\land c$ $a\land b\land c$
@@ 3.1 Static rationality
In the next step, we will show that it is possible to characterize linear probabilistic aggregation rules for a very large class of agendas – all non-nested agendas. In the setting of probabilistic opinion pooling, the agenda has been traditionally assumed to be a $\sigma$-algebra (i.e. closed under complementation and countable union of events). Dropping the assumption of a $\sigma$-algebra, [9] consider more general agendas. The constraint of nestedness implies that formulas $\phi_{1},...,\phi_{r}\in X$ are logically interconnected by material implication which corresponds to the subset-relations in a nested agenda $X$ in the framework of [9]. Intuitively, non-nestedness means that there is no subagenda in $X$ that comprises only propositions that are logically interrelated by material implications and that are closed under negation. Thus, the propositions under consideration can be probabilistically dependent (correlated) withour being logically interrelated by material implication.
@@
Let $S=\{0,1\}^{|A|}$ be the set of all truth assignments to the finite set of atomic propositions $A$. For every formula $\phi\in\mathcal{L}$ define $[\phi]=\{v\in S\mid v\text{ satisfies }\phi\}.$ It is easy to show that the mapping $I:\mathcal{L}\to 2^{S},\ \phi\mapsto[\phi]$ is bijective. Since $I$ preserves the Boolean operations, i.e.,
@@
Individual $a$ $b$ $c$ $a\rightarrow b$ $a\land b$ $a\land c$ $b\land c$ $a\land b\land c$
@@
$F(J_{1},J_{2},J_{3})$ 0.7 0.5667 0.4 0.7667 0.4667 0.3 0.2333 0.1667
Table 2 demonstrates the resulting collective judgement in the running example using the linear aggregation rule. Note that the conditions of Theorem 3.1 (e.g. non-nestedness, number of contingent elements) are satisfied in our example. Intuitively, Theorem 3.1 does not exclude the possibility of non-dictatorial and probabilistically rational aggregation of individual attitudes in our framework. Non-dictatorship means there is there is no individual $i\in N$ such that, for every profile $(J_{1},\ldots,J_{n})$ in the domain of $F$ and every formula $\phi\in X$, the collective judgement for $\phi$ is given by $F(J_{1}(\phi),\ldots,J_{n}(\phi))=J_{i}(\phi)$. Linearity encompasses both the case of dictatorship (if one weight equals 1 and the others 0) and the possibility of non-dictatorial aggregation (if the weights are chosen otherwise). The possibility of non-dictatorial aggregation of individual attitudes on logically interrelated issues arises because the domain of the probabilistic aggregation rule $F$ is extended in our model allowing for probabilistic individual attitudes, as compared to impossibility results from the standard framework of judgement aggregation (see e.g. [16, 3, 19]).
@@ 3.2 Dynamic Rationality
The requirement of probabilistic dynamic rationality is a plausible constraint, since it demands that the individual and collective probabilistic judgements are updated after learning the truth of $\phi$ in such a manner so that it makes no difference whether they are updated before aggregation or after aggregation process. Thus, it intuitively corresponds to the requirement of external Bayesianity from the literature on probability aggregation [11] and to the condition of dynamic rationality in the judgement aggregation framework of [10]. The following result illustrates that the possibility of dynamically rational aggregation with a $\Phi$-preserving update operator on the restricted domain $\mathcal{D}_{\Phi}$.
@@
Individual $a$ $b$ $c$ $a\rightarrow b$ $a\land b$ $a\land c$ $b\land c$ $a\land b\land c$
@@
Table 4: Updated probabilistic judgments after learning $\neg c$, following the update on $a$, with values rounded to four decimals. The probabilities for $b$ and $a\rightarrow b$ are computed as $J^{a,\neg c}(b)=J^{a,\neg c}(a\rightarrow b)=\frac{J^{a}(b\land\neg c)}{J^{a}(\neg c)}=\frac{J^{a}(b)-J^{a}(b\land c)}{J^{a}(\neg c)}$, and for $a\land b$ as $J^{a,\neg c}(a\land b)=\frac{J^{a}((a\land b)\land\neg c)}{J^{a}(\neg c)}=\frac{J^{a}(a\land b)-J^{a}(a\land b\land c)}{J^{a}(\neg c)}$, using the value $J^{a}(\neg c)=0.5714$. Here, $F(*)$ abbreviates $F(U(J_{1},a,\neg c),U(J_{2},a,\neg c),U(J_{3},a,\neg c))$.
Individual $\ a$ $b$ $c$ $a\rightarrow b$ $a\land b$ $a\land c$ $b\land c$ $a\land b\land c$
@@
A key aspect of our model is the restricted domain $\mathcal{D}_{\Phi}$, which is justified by the notion of fair learning. Here, learning is deemed fair when group members can only update on propositions in the common ground $\Phi\subseteq X$, where all members have equal degrees of confidence (i.e., $\forall\phi\in\Phi,\ \forall i,j\in N:J_{i}(\phi)=J_{j}(\phi)$). Under fair learning, new information about any $\phi\in\Phi$ benefits all members equally, thereby preserving dynamic rationality.
Our result points out that impossibility theorems in dynamically rational judgment aggregation such as [10] arise when fair learning is absent—namely, when members begin with differing degrees of confidence. Our model not only supports fair learning but also underscores its necessity for maintaining rationality in dynamic collective decision-making.
Furthermore, while previous research (e.g., [11]) has shown that geometric pooling functions satisfy external Bayesianity and unanimity preservation, and [13] demonstrated that the linear opinion pool fails external Bayesianity unless it is dictatorial, our framework adopts a different perspective. In Theorem 3.2 we show that within our propositional probability logic framework, linear aggregation can be dynamically rational under suitable domain restrictions. This result does not conflict with earlier findings but rather highlights that linear averaging, when combined with our approach to belief updating, yields a dynamically rational aggregation rule—an insight that opens promising new avenues for research and applications in dynamic and sequential collective decision-making.
@@ 4 Concluding Remarks
Building on the broad implications of Theorem 3.2, we conclude that our framework—rooted in propositional probability logic—provides a generalization of classical judgment and preference aggregation approaches, thereby offering new insights into dynamic and sequential collective decision-making. We have demonstrated that, under the constraints of consensus compatibility and independence on a non-nested agenda, any aggregation rule must be linear. More importantly, by restricting the domain to a common ground (i.e., enforcing fair learning), dynamic rationality is achieved: Bayesian updating, whether performed before or after aggregation, yields the same collective judgment. A notable contribution of our dynamically rational model is its emphasis on sequential decision-making. The property of common ground preservation across successive updates is crucial in multi-stage decision-making contexts, where new evidence on issues under consideration emerges progressively.
Our positive result highlights the crucial role of fair learning in ensuring dynamic consistency in collective decision-making. While our analysis leaves open normative considerations on the choice of aggregation weights (which may lead to dictatorial outcomes in extreme cases), it clearly shows that non-dictatorial and dynamically rational aggregation is attainable under appropriate conditions.
@@ 4.0.1 Acknowledgements
Polina Gordienko gratefully acknowledges the support of the Friedrich-Ebert-Stiftung Academic Foundation.

### v2

@@ 1 Introduction
In contrast, this paper advocates a shift toward a dynamic theory of collective decision-making in which both individual and collective beliefs update rationally in response to new information (see [5]). Motivated by the first impossibility result for dynamically rational judgment aggregation [10], our main contribution is a possibility result for dynamically rational collective decision-making on logically interconnected propositions. We introduce a framework of probability aggregation based on propositional probability logic that generalizes traditional approaches from both judgment and preference aggregation. Building on prior work in probability aggregation [17, 1, 2, 12, 18, 7, 9], our first theorem establishes that any consensus-compatible and independent aggregation rule on a non-nested agenda must be linear. Furthermore, we show that dynamically rational linear aggregation is attainable when the domain of the aggregation function is restricted to profiles in which all group members share the same probabilistic judgments for a specified subset of propositions—what we term the common ground (here understood in line with a broader notion of consensus) [^1] Throughout this paper, “consensus” is used in two senses: (i) as a broad notion of mutual agreement—here, synonymous with our technical term “common ground” and (ii) as “consensus compatibility,” the requirement that if all individuals assign probability 1 to a proposition, then the collective judgment must also assign 1.. This common ground represents the set of propositions on which all individuals already agree, and by ensuring that new information is limited to this shared foundation—a process we call fair learning—all members benefit equally, preventing dynamically irrational outcomes arising from unequal starting beliefs. Importantly, we demonstrate that fair learning not only achieves dynamic rationality but also preserves the common ground across sequential updates.
@@ 2.1 Propositional probability logic
Let $\mathcal{L}$ be the set of all logically non-equivalent [^2] Logical equivalence (i.e., $\phi\equiv\psi$ if and only if $\models\phi\leftrightarrow\psi$) is an equivalence relation. We select one finitary representative from each equivalence class so that no two distinct formulas in $\mathcal{L}$ are equivalent. Since $A$ is finite, and we consider only finitary formulas, $\mathcal{L}$ is finite. formulas based on a propositional logic with a finite set $A$ of atomic formulas. The variable $\phi$ is then a generic variable ranging over all well-formed formulas in $\mathcal{L}$, i.e., any expression that can be constructed from the atomic elements in $A$ using the logical connectives $\neg,\land,\lor,\rightarrow,\leftrightarrow$.
@@
  • If $\phi_{1},\phi_{2}$$\in\mathcal{L}$ such that $\models\neg(\phi_{i}\land\phi_{j})$ for all $i\neq j\in\mathbb{N}$, then
    $P(\bigvee_{i=1}^{\infty}\phi_{i})=\sum_{i=1}^{\infty}P(\phi_{i}).$
@@ 2.2 The framework of probability aggregation
The probabilistic judgement $J$ is a function $J\colon X\rightarrow[0,1]$. The judgement $J$ is called probabilistically rational if there exists a propositional probability $P^{*}\colon\mathcal{L}\rightarrow[0,1]$ on the whole set $\mathcal{L}$ such that $\forall\ \phi\in X\colon J(\phi)=P^{*}(\phi)$. A probabilistic profile on an agenda $X$ is a tuple $(J_{1},...,J_{n})$, where each of $J_{1},...,J_{n}$ is a probabilistically rational probabilistic judgement. The probabilistic aggregation rule is a function $F\colon J(X)^{n}\rightarrow J(X),$ where $J(X)$ is the collective set of all probabilistically rational probabilistic judgements on the agenda $X$.
@@ 2.3 A Running Example: A Political Scenario
Population ageing is straining the healthcare and pension systems.
@@
The current tax revenue is sufficient to meet the fiscal demands imposed by population ageing.
@@
Table 1: A sample probabilistic profile on the agenda $X=\pm\{a,b,c,a\rightarrow b,a\land b,a\land c,b\land c,a\land b\land c\}$, where each individual assigns a degree of confidence (ranging from 0 to 1) to the propositions. Each entry is chosen so as to align with the relationships dictated by logical connectives, ensuring that bounds for probability assignments are satisfied. Note that since $X$ is $\wedge$-stable, this agenda uniquely determines a single joint distribution over $\{a,b,c\}$.
Individual aa\ \ \ \ bb\ \ \ \ cc\ \ \ \ $a\rightarrow b\$$a\land b\$$a\land c\$$b\land c\$$a\land b\land c\$
@@ 3.1 Static rationality
In the next step, we will show that it is possible to characterize linear probabilistic aggregation rules for a very large class of agendas – all non-nested agendas. In the setting of probabilistic opinion pooling, the agenda has been traditionally assumed to be a $\sigma$-algebra (i.e. closed under complementation and countable union of events). Dropping the assumption of a $\sigma$-algebra, [9] consider more general agendas. The constraint of nestedness implies that formulas $\phi_{1},...,\phi_{r}\in X$ are logically interconnected by material implication which corresponds to the subset-relations in a nested agenda $X$ in the framework of [9]. Intuitively, non-nestedness means that there is no subagenda in $X$ that comprises only propositions that are logically interrelated by material implications and that are closed under negation. Thus, the propositions under consideration can be probabilistically dependent (correlated) without being logically interrelated by material implication.
@@
[…]
@@
Individual aa\ \ \ \ bb\ \ \ \ cc\ \ \ \ $a\rightarrow b\$$a\land b\$$a\land c\$$b\land c\$$a\land b\land c\$
@@
$F(J_{1},J_{2},J_{3})\$0.7 0.5667 0.4 0.7667 0.4667 0.3 0.2333 0.1667
Table 2 demonstrates the resulting collective judgement in the running example using the linear aggregation rule. Note that the conditions of Theorem 3.1 (e.g. non-nestedness, number of contingent elements) are satisfied in our example. Intuitively, Theorem 3.1 does not exclude the possibility of non-dictatorial and probabilistically rational aggregation of individual attitudes in our framework. Non-dictatorship means there is no individual $i\in N$ such that, for every profile $(J_{1},\ldots,J_{n})$ in the domain of $F$ and every formula $\phi\in X$, the collective judgement for $\phi$ is given by $F(J_{1}(\phi),\ldots,J_{n}(\phi))=J_{i}(\phi)$. Linearity encompasses both the case of dictatorship (if one weight equals 1 and the others 0) and the possibility of non-dictatorial aggregation (if the weights are chosen otherwise). The possibility of non-dictatorial aggregation of individual attitudes on logically interrelated issues arises because the domain of the probabilistic aggregation rule $F$ is extended in our model allowing for probabilistic individual attitudes, as compared to impossibility results from the standard framework of judgement aggregation (see e.g. [16, 3, 19]).
@@ 3.2 Dynamic Rationality
The requirement of probabilistic dynamic rationality is a plausible constraint, since it demands that the individual and collective probabilistic judgements are updated after learning the truth of $\phi$ in such a manner that it makes no difference whether they are updated before aggregation or after aggregation process. Thus, it intuitively corresponds to the requirement of external Bayesianity from the literature on probability aggregation [11] and to the condition of dynamic rationality in the judgement aggregation framework of [10]. The following result illustrates that the possibility of dynamically rational aggregation with a $\Phi$-preserving update operator on the restricted domain $\mathcal{D}_{\Phi}$.
@@
Individual aa\ bb\ \ \ \ \ cc\ \ \ \ \ $a\rightarrow b\ \$$a\land b\ \$$a\land c\ \$$b\land c\ \$$a\land b\land c\ \$
@@
Table 4: Updated probabilistic judgments after learning $\neg c$, following the update on $a$, with values rounded to four decimals. The probabilities for $b$ and $a\rightarrow b$ are computed as $J^{a,\neg c}(b)=J^{a,\neg c}(a\rightarrow b)=\frac{J^{a}(b\land\neg c)}{J^{a}(\neg c)}=\frac{J^{a}(b)-J^{a}(b\land c)}{J^{a}(\neg c)}$, and for $a\land b$ as $J^{a,\neg c}(a\land b)=\frac{J^{a}((a\land b)\land\neg c)}{J^{a}(\neg c)}=\frac{J^{a}(a\land b)-J^{a}(a\land b\land c)}{J^{a}(\neg c)}$, using the value $J^{a}(\neg c)=0.5714$. Here, F()F(*) abbreviates $F(U(J_{1},a,\neg c),U(J_{2},a,\neg c),U(J_{3},a,\neg c))$.
Individual a\ a\ \ bb\ \ \ \ \ cc\ \ \ $a\rightarrow b\ \$$a\land b\ \$$a\land c\ \$$b\land c\ \$$a\land b\land c\ \$
@@
A key aspect of our model is the restricted domain $\mathcal{D}_{\Phi}$, which is justified by the notion of fair learning. Here, learning is deemed fair when group members can only update on propositions in the common ground $\Phi\subseteq X$, where all members have equal degrees of confidence (i.e., $\forall\phi\in\Phi,\ \forall i,j\in N:J_{i}(\phi)=J_{j}(\phi)$). Under fair learning, any Bayesian update on $\phi\in\Phi$ produces the same posterior shift in every individual’s beliefs, thereby preserving dynamic rationality. Our epistemic notion of fairness is grounded in equal information processing, guaranteeing that no agent is advantaged or disadvantaged by newly received data.
It has been established in [10] that under standard aggregation axioms (e.g. universal domain, systematicity) no judgment aggregation rule can be dynamically rational with respect to any revision operator satisfying basic conditions on revision, if the propositions in the agenda are non-trivially interrelated. By contrast, when we recast the same agenda in our probabilistic framework – replacing “yes/no” verdicts with probability assignments over the logically interconnected propositions – and impose a common-ground domain restriction – weighted linear pooling both preserves collective coherence and commutes with Bayesian updating.
Thus, our result points out that impossibility theorems in dynamically rational judgment aggregation such as [10] arise when fair learning is absent – namely, when members begin with differing degrees of confidence. Our model not only supports fair learning but also underscores its necessity for maintaining rationality in dynamic collective decision-making.
Furthermore, while previous research (e.g., [11]) has shown that geometric pooling functions satisfy external Bayesianity and unanimity preservation, and [13] demonstrated that the linear opinion pool fails external Bayesianity unless it is dictatorial, our framework adopts a different perspective. In Theorem 3.2 we show that within our propositional probability logic framework, linear aggregation can be dynamically rational under suitable domain restrictions. This result does not conflict with earlier findings but rather highlights that linear averaging, when combined with our approach to belief updating, yields a dynamically rational aggregation rule – an insight that opens promising new avenues for research and applications in dynamic and sequential collective decision-making.
@@ 4 Concluding Remarks
Building on the implications of Theorem 3.2, we conclude that our framework generalizes classical judgment and preference aggregation approaches, offering new insights into dynamic and sequential collective decision-making. We demonstrated that under consensus compatibility and independence on a non-nested agenda, any aggregation rule must be linear. More importantly, when the domain is restricted to a common ground (i.e., under fair learning), dynamic rationality is achieved: Bayesian updating, whether performed before or after aggregation, yields the same collective judgment. A key contribution of our model is its emphasis on sequential decision-making, where common ground preservation across updates enables consistent multi-stage deliberation. Our result highlights the central role of fair learning in ensuring dynamic consistency and avoiding the impossibilities found in more general domains. While we do not directly address the normative choice of aggregation weights, our findings show that non-dictatorial and dynamically rational aggregation is possible under appropriate conditions.
@@ 4.0.1 Acknowledgements
Polina Gordienko gratefully acknowledges the support of the Friedrich-Ebert-Stiftung Academic Foundation. We would like to thank three anonymous referees for their helpful comments.
4.0.2 Disclosure of Interests
The authors have no competing interests to declare that are relevant to the content of this article.